OpenStudy (anonymous):

For anyone who knows how to solve a system of equations in order to find a quadratic function.... (0,15) (2,5) (3,6). Using elimination.

OpenStudy (anonymous):

\[f(x)=ax^2+bx+c\] \[f(0)=c=15\] so \[f(x)=ax^2+bx+15\] is a start

OpenStudy (anonymous):

\[f(2)=a\times (2)^2+\times 2+15=5\\ 4a+2b+15=5\\ 4a+2b=-10\] is one equation

OpenStudy (anonymous):

repeat with \(f(3)=6\) and you will have another equation in \(a\) and \(b\) then you can solve

OpenStudy (anonymous):

I got that far and have these three equations. What do I do with the 15? 15 = C 5 = 4a + 2b + C 6 = 9a + 3b + C

OpenStudy (anonymous):

\(c=15\)

OpenStudy (anonymous):

i.e. replace \(c\) by \(15\) in your two equations

OpenStudy (anonymous):

so I just get rid of C from the equations by subtracting

OpenStudy (anonymous):

\[5 = 4a + 2b + C\\ 6 = 9a + 3b + C\] \[5 = 4a + 2b + 15\\ 6 = 9a + 3b + 15\]

OpenStudy (anonymous):

yes, by subtracting \(15\) to get \[-10=4a+2b\\ -9=9a+3b\]

OpenStudy (anonymous):

Then elimination?

OpenStudy (anonymous):

-30 = 12a + 6b -18 = 18a + 6b

OpenStudy (anonymous):

Does A = 2?

OpenStudy (anonymous):

B = -9? btw thanks for all the help

OpenStudy (anonymous):

hello?